Circular No. 5745

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET SHOEMAKER-LEVY (1993e) Perturbed ephemeris, based on the elements on IAUC 5744:

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Feb. 22 12^{h}39.^{m}50 − 5^{o}15.^{′}5 4.562 5.379 142.^{o}3 6.^{o}5 13.6
Mar. 4 12 35.92 − 4 56.6 4.471 5.376 153.3  4.8 13.6
14 12 31.60 − 4 32.4 4.408 5.373 164.4  2.8 13.5
24 12 26.83 − 4 04.6 4.375 5.369 175.6  0.8 13.5
Apr. 3 12 21.94 − 3 35.3 4.372 5.366 173.0  1.3 13.5
13 12 17.28 − 3 06.8 4.400 5.363 162.0  3.3 13.5
23 12 13.16 − 2 40.9 4.457 5.360 151.1  5.2 13.5
May 3 12 09.84 − 2 19.7 4.540 5.357 140.5  6.9 13.6
13 12 07.51 − 2 04.3 4.645 5.354 130.2  8.3 13.6
23 12 06.28 − 1 55.6 4.769 5.350 120.2  9.4 13.7
June 2 12 06.20 − 1 54.0 4.906 5.347 110.6 10.2 13.7

The reported positions of the comet refer to the center of the train (or bar). As noted on IAUC 5730, the brightest nuclei are displaced to the west-southwest. The Hawaiian observations seem to show a very close pair of nuclei there, and it does not necessarily follow that this is the comet’s center of mass. J. V. Scotti, Lunar and Planetary Laboratory, reports that Spacewatch observations on Mar. 28.24 UT showed 11 nuclei, essentially along p.a. 76^{o}-256^{o}. On the western side of the center nuclei were measured at distances 25′′.4, 20′′.9, 16′′.7, 13′′.4, 8′′.8 (the brightest, estimated at m_{2} = 20.2) and 1′′.4 from the center. On the eastern side the nuclei were at distances 3′′.2, 9′′.4, 13′′.7, 19′′.7 and 27′′.0. A tail extended ∼ 1′ from the train as a whole, and a brighter component of the tail extended 1′.21 in p.a. 285^{o} from the brightest nucleus. A dust trail extended at least 4′.37 in p.a. 258^{o} from the midpoint of the train. A second dust trail extended 3′.96 in p.a. 72^{o}. The brighter nuclei were also detected in the Mar. 26.30 record (cf. IAUC 5725); these were measured at 20′′.6 and 9′′.7 (the brightest) and 0′′.7 to the west of center and at 11′′.7 and 20′′.7 to the east. E. M. Shoemaker reports that three of the nuclei are also detectable in enlargements of the discovery films: the Mar. 24.43 exposure shows enhancements at 19′′.1, 12′′.6 (the brightest) and 5′′.5 on the western side of center.

Total visual magnitude estimates: Mar. 28.39 UT, 13.3 (A. Hale, Las Cruces, NM, 0.41-m reflector); 30.39, 12.1 (D. H. Levy, Tucson, AZ, 0.40-m reflector); 31.38, 12.3 (Levy); 31.40, 13.4 (Hale).

1993 April 3 Brian G. Marsden